OpenStudy (anonymous):

pls help with this, if r=d/2 find the area of the bigger semi-circle, subtract the area of the smaller semi-circle from it to get the area of the shaded region. 5in 10 in-shaded area

OpenStudy (saifoo.khan):

big radius : 10in small radius : 5 in??

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

58.9in^2

OpenStudy (anonymous):

i need to see work because i had that and my teacher said its wrong

OpenStudy (anonymous):

a = pi(r^2) a = pi(5/2)^2 a = 25/4pi area larger circle is a = pi(r^2) a = pi(10/2)^2 a = pi(5^2) a = 25pi larger - smaller = 25pi - 25/4pi = 100/4pi - 25pi = 75pi/4 is wat i got

OpenStudy (saifoo.khan):

are the given measurements in Radius or Diamerter?

OpenStudy (dumbcow):

shenita area of semi-circle is (pi*r^2)/2 take half the area, you were squaring (r/2) area of big: (pi*10^2)/2 = 100pi/2 = 50pi area of small: (pi*5^2)/2 = 25pi/2 = 12.5pi subtract 50pi - 12.5pi = 37.5pi = 117.81

OpenStudy (anonymous):

ook i see thanx